payton:

What is the solution to the system of equations? y = 1/3x - 10 2x + y = 4 (–8, 6) (–6, 16) (6, –8) (16, –6)

lowkey:

Plug in the first formula unto the second formula in place of the y variable

lowkey:

2x+1/3x-10=4 7/3x-10=4 7/3x=14 x=6

lowkey:

What answer choice gives you 6 as an x value

payton:

only option would be c

lowkey:

Yes good job

payton:

couldnt have done it without you

lowkey:

you're welcome! I'm able to help you whenever i'm online

payton:

thx

lowkey:

No prob